OpenStudy (anonymous):

A cheetah can go from a state of resting to running at 20 m/s in just 2 seconds. What is the cheetah's average acceleration? A. 40 m/s B. 10 m/s² C. 10 m/s D. 40 m/s²

OpenStudy (anonymous):

a = v/t a = 20/2 a = 10ms-2

OpenStudy (anonymous):

yes, a =(end vel - start vel)/(time change) = 20/2